OpenStudy (anonymous):

The hypotenuse of a right triangle is 7 inches, and one of the legs is inches. Find the length of the other leg.

OpenStudy (rmrjr22):

a^2+b^2=c^2

OpenStudy (rmrjr22):

c^2 = hypotnuse

OpenStudy (rmrjr22):

ur missing a variable

OpenStudy (anonymous):

yeah where it says and one of the legs is sqrt13 inches

OpenStudy (anonymous):

 The location of the sqrt(13) and the unknown can be on opposite sides (a or b) but the hypotenuse is the longest side so it will be c in a^2+b^2=c^2 sqrt(13)+b^2=7^2 13+b^2=49 solve for c (be careful of the square root)

OpenStudy (rmrjr22):

\[\sqrt{13}^2+b^2 = 7^2\]

OpenStudy (rmrjr22):

find b

OpenStudy (anonymous):

thanks for the help I got the answer